OpenStudy (xapproachesinfinity):

check the convergence of the following: \[\sum_{n=1}^{\infty}\frac{1}{n^{\frac{n+1}{n}}}\] I used log to show it is divergent \[\ln(\frac{1}{L})=\sum \ln n + \sum \frac{\ln n}{n}\]

OpenStudy (xapproachesinfinity):

how can i do it with comparison test

OpenStudy (xapproachesinfinity):

however, I'm not sure if my line of reasoning was correct considering the use of logs

OpenStudy (xapproachesinfinity):

@amistre64

OpenStudy (xapproachesinfinity):

@Kainui

OpenStudy (amistre64):

not really a strong point of mine, ganesha seems to work these like they were 3rd grade arithmetic

OpenStudy (xapproachesinfinity):

some i struggle with comparison test some cases it is obvious what series you need to compare, but some cases it is not so clear

OpenStudy (xapproachesinfinity):

somehow*

OpenStudy (amistre64):

@rational you got any insights?

OpenStudy (xapproachesinfinity):

yeah ganesh is good with these stuff lol how about my approach any missteps?

OpenStudy (xapproachesinfinity):

can we compare that with n!?

OpenStudy (xapproachesinfinity):

the question asks for comparison test! tried other tests, they fail inclusive!

OpenStudy (rational):

tried root test yet ?

OpenStudy (xapproachesinfinity):

yes failed! I my attempt was correct

OpenStudy (xapproachesinfinity):

if*

OpenStudy (anonymous):

stumped i bet $50 it diverges since the exponent goes to one, but i'll be damned if i see what to compare it to

OpenStudy (rational):

try comparing with 1/(n*logn)

OpenStudy (xapproachesinfinity):

it is divergent 100%, problem is how to prove it is

OpenStudy (xapproachesinfinity):

so 1/nlogn is less than that expression?

OpenStudy (xapproachesinfinity):

comparison test give me hard time lol how did you figure it out that would be good for to compare it with

OpenStudy (rational):

because SUM 1/(n*ln(n)) is special and very famous it is the middle ground between convergence and divergence of p-series since we had a feeling that the given series diverges, and it is kinda similar to harmonic series...

OpenStudy (xapproachesinfinity):

but need to show somehow that it is less than the given series to use comparison test

OpenStudy (anonymous):

i didn't know there was a middle ground, but i knew it diverged but so does \[\sum \frac{1}{n\log(n)\log(\log(n))}\] if i remember correctly

OpenStudy (xapproachesinfinity):

yeah that diverges by integral test

OpenStudy (rational):

1/n diverges `1/(n*ln(n)) diverges` 1/n^2 converges

OpenStudy (xapproachesinfinity):

hmm i see what you mean saw something like this on mit open course

OpenStudy (rational):

any order greater than O(n*ln(n)) in the denominator makes the p-series converge

OpenStudy (rational):

yeah i remember it from there too

OpenStudy (anonymous):

oh what a good idea !

OpenStudy (anonymous):

show \[n^{\frac{n+1}{n}}<n\log(n)\] cool

OpenStudy (xapproachesinfinity):

hmm good question induction

OpenStudy (anonymous):

take the log lol

OpenStudy (rational):

or to make ur life more simple, try limit comparison test

OpenStudy (xapproachesinfinity):

oh yeah that's a good idea

OpenStudy (xapproachesinfinity):

limit will be ugly hehe

OpenStudy (xapproachesinfinity):

\[\lim \frac{1}{n\log( n) n^{\frac{n+1}{n}}}\]

OpenStudy (xapproachesinfinity):

what would that limit go to?

OpenStudy (xapproachesinfinity):

let me try wolfram no will to do it hehe

OpenStudy (xapproachesinfinity):

goes to zero no good lol

OpenStudy (rational):

yeah try comparison oly as satallite said

OpenStudy (xapproachesinfinity):

need to show that inequality i will try to see if that holds for some n or any n

OpenStudy (xapproachesinfinity):

not good for any n

OpenStudy (xapproachesinfinity):

n>=5 seems where it is true

OpenStudy (rational):

try this For all positive \(n\), we have \(\large \sqrt[n]{n} \lt 2\)

OpenStudy (xapproachesinfinity):

what is the goal?

OpenStudy (rational):

\[\large {\sqrt[n]{n} \lt 2 \implies \frac{1}{\sqrt[n]{n}}\gt \frac{1}{2} \implies \frac{1}{n*\sqrt[n]{n}}\gt \frac{1}{2n} }\]

OpenStudy (xapproachesinfinity):

i see 1/2n diverges

OpenStudy (rational):

that will do yeah

OpenStudy (xapproachesinfinity):

that's pretty good actually :)

OpenStudy (xapproachesinfinity):

great job man :)

OpenStudy (rational):

np :) is it easy to show that \(\sqrt[n]{n}\) is less than 2

OpenStudy (xapproachesinfinity):

hmm let me do that lol direct proof

OpenStudy (xapproachesinfinity):

i took it to \[\ln \frac{n}{2^n}<0\] am i deviating

OpenStudy (xapproachesinfinity):

since n/2^n <1 for any n then ln(n/2^n)<0 have to start from here and go backwards

OpenStudy (xapproachesinfinity):

sounds good?

OpenStudy (rational):

thats pretty clever!

OpenStudy (xapproachesinfinity):

:)

OpenStudy (xapproachesinfinity):

then we are done :) thanks a lot feel good this is solved lol

OpenStudy (rational):

:D

OpenStudy (rational):

that ln(1/L) thingy really worked ?

OpenStudy (rational):

im referring to the main q http://gyazo.com/a9621914c1299a48d614cbed5c2888e8

OpenStudy (xapproachesinfinity):

will the left hand side is divergent but the reasoning lacks something hehe

OpenStudy (rational):

Okay that looked bit unorthodox haha! but its a good idea

OpenStudy (xapproachesinfinity):

just tried a bunch of things lol no harm in exploring:)